Since a coin is weighted, tails is more likey. When you test the coin by flipping it 10 times, you observe that tails comes up 9 times. How likely is it that such extreme behavior would occur in a fair coin?(In other words, assuming that the coin is fair, what is the probability of getting tails 9 or 10 times?)
Answer #1

Binomial distribution: P(x) = nCx px qn-x

P(tails), p = 0.5

q = 1 - p = 0.5

number of tosses, n = 10

Probability of getting 9 or 10 tails in fair coin = P(9) + P(10)

= 10C9 x 0.59 x 0.5 + 0.510

= 0.0107